Define uniform convergence of a sequence of functions. Discuss the uniform convergence of the sequence nx^2+1/nx+1 the interval  [1. 2]

1
Expert's answer
2021-05-26T03:31:59-0400

In a hyper-real context, uniform convergence allows for a simpler description. So, if is infinitely close to for all x in the space of and all infinite m, a sequence converges to f uniformly.(For a related concept of uniform continuity, see micro continuity)

"f(x)=\frac{nx^2+1}{nx+1}"

"\lim\limits_{x \to \infin}f_n(x)=\lim\limits_{x \to \infin}[\frac{nx^2+1}{nx+1}]=\lim\limits_{x \to \infin}[\frac{\frac{n}{n}x^2+\frac{1}{n}}{\frac{nx}{n}+\frac{1}{n}}]=\lim\limits_{x \to \infin}[\frac{x^2+\frac{1}{n}}{x+\frac{1}{n}}]=\lim\limits_{x \to \infin}[\frac{x^2+\frac{1}{\infin}}{x+\frac{1}{\infin}}]= x"

When x=0

"f_n(x)=0"

Therefore, "f(x)={x \space if \space x \not =0\brack 0 \space if \space x=0}"

"\lim\limits_{x \to \infin}f_n(x)= f(x)" for all x.

To prove that the above convergence is not pointwise

Suppose that "f_n \to f" uniformly

Therefore, "\exist \varepsilon >0, \exist x>[1,2], \forall N \epsilon \N, \exist n> N"

"|f_n(x)-f(x)|< \varepsilon"

Now , taking "\varepsilon = \frac{1}{10}, x=\frac{1}{\sqrt{n}}" and n=N+1

Therefore, "|\frac{nx^2+1}{nx+1}-x|=\frac{1-x}{nx+1}>\varepsilon=\frac{1}{10}"

Which is a contradiction to the fact that "f_n(x)" converges uniformly

Hence , "f_n(x)" does not converge to f uniformly.